OpenStudy (anonymous):

what is the factor of 3w2+5w+2

jhonyy9 (jhonyy9):

you can rewrite this quadratic in the form of 3w^2 +3w +2w +2 yes ? so now from first two and from 3rd and 4rd what you can factoriz out ?

jhonyy9 (jhonyy9):

are you here ?

jhonyy9 (jhonyy9):

courage this is easy

OpenStudy (anonymous):

yea

jhonyy9 (jhonyy9):

so from first two term you can factorizing 3w and what will get ?

jhonyy9 (jhonyy9):

3w(w+1) yes and from 3rd and 4rd you can factorizing 2 and will get 2(w+1) yes ?

jhonyy9 (jhonyy9):

so than now you see what can being the factor of quadratic ok ?

OpenStudy (anonymous):

ok so whats is the two terms

jhonyy9 (jhonyy9):

3w(w+1)+2(w+1)=(3w+2)(w+1) do you understand it now sure ?

OpenStudy (anonymous):

what about 2f2+13f+11

jhonyy9 (jhonyy9):

so there are 3w(w+1)+2(w+1)= what can you here now factorizing out ?

OpenStudy (anonymous):

what is the factor of 2f2+13f+11

OpenStudy (anonymous):

what is the two tedms

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

good luck was my pleasure

jhonyy9 (jhonyy9):

rewrite it 2f^2 +2f +11f +11 = how you continue ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

whats the next thing

jhonyy9 (jhonyy9):

so factoriz out 2f and from 3rd and 4rd 11 so what will get

OpenStudy (anonymous):

what is the two terms

jhonyy9 (jhonyy9):

is the same life what i have solved above

jhonyy9 (jhonyy9):

2f(f+1)+11(f+1) and now continue exactly like the first exersice ok ?

OpenStudy (anonymous):

what about 8t2-14t+3